OpenStudy (shubhamsrg):

0< a,b,c <1 if a+b+c=2 prove that (abc)/( (1-a)(1-b)(1-c) ) >=8

OpenStudy (anonymous):

I'm not sure if this is helpful, but since \(a,b,c\in(0,1)\), you could rewrite \[\frac{abc}{(1-a)(1-b)(1-c)}\] as the product of infinite series: \[\sum_{n=1}^\infty a^n\cdot \sum_{n=1}^\infty b^n \cdot \sum_{n=1}^\infty c^n\] But like I said, I don't know where this would take you, nor do I see how to use the fact that \(a+b+c=2\).

OpenStudy (anonymous):

 Now do this, (1-a) + (1-b) + (1-c) = 1 and 0< (1-a) <1, 0<(1-b)<1 and 0<(1-c)<1  Almost done... just divide and notice that the inequality sign will change...